Which of the following sentences is a proposition whose truth value is T?


Select one:

a. x + 2 = 11

b. x + y = y + x for every pair of real numbers x and y

c. Answer this question carefully.

d. Manila is in Cagayan.




1
Expert's answer
2022-02-06T17:00:08-0500

"a)"

"x + 2 = 11."

The given statement can neither be true or false. It depends on the value of x. Hence, it is not a proposition.


"b)"

x + y = y + x for every pair of real numbers x and y

This is a proposition whose truth value is "T."

"c)"

 Answer this question carefully.

Since " Answer this question carefully" is not a statement that can be either true or false, it is not a proposition.


"d)"

Manila is in Cagayan.

This statement is a proposition whose truth value is "F" .

Therefore, the correct choice is b.